# Reviewing Uses of Regulatory Compliance Monitoring

### v4

@@ 3. RQ3:
This paper is based on a previous study, where we reviewed conformance checking uses for regulatory compliance monitoring purposes [klessascheckReviewingConformance2024]. Extending our focus, we now additionally investigate usages of further techniques that exist in the research space of regulatory compliance and BPM in our review, and take additional aspects such as the origin and availability of process execution data used in the investigated studies into account. We also describe our research method in more detail and, in light of our broadened scope, extend the results and discussion provided herein.
@@ 2.2 Related Work
Building on this framework, Rinderle-Ma et al. [rinderle-maPredictiveComplianceMonitoring2023] classify predictive compliance monitoring approaches along compliance monitoring requirements and capabilities. For doing so, they extend the requirements formulated by Ly et al. [lyComplianceMonitoringBusiness2015] with e.g. additional user requirements or data requirements specific to predictive process compliance monitoring. Similar to Ly et al. [lyComplianceMonitoringBusiness2015], they assess in how far existing literature covers the functionality requirements they elicited, and derive research directions from this assessment. Requirements towards compliance monitoring approaches are elicited from real-world case studies and literature in both studies; however, concrete applications of these approaches are not characterized and investigated for their particular properties. While the requirements are useful for describing the capabilities of compliance monitoring approaches, they are insufficient to characterize how these approaches are applied in a way that goes beyond their ability to meet e.g. modelling, execution, and user requirements. Instead, we believe more emphasis should be placed on how prescriptive models come to be, to what end uses of approaches are reported, and in what ways the results are represented (see also e.g. Rehse et al. [rehse2022process]), in particular for regulatory compliance.
@@ 3 Research Method
In Search B, we explicitly search for studies that position themselves as applying techniques besides conformance checking (meaning, other compliance checking techniques) for regulatory compliance monitoring. In line with Search A, we include terms commonly associated with regulations. The complete search term ((“compliance checking” OR “compliance monitoring”) AND (“law” OR “guideline” OR “manual” OR “reference” OR “quality” OR “compliance” OR “regulation”)) is applied to the title, abstract, and keyword search, where applicable. [^1] Note that for the search we use compliance monitoring and compliance checking interchangeably for any study that mentions applications of compliance monitoring or checking techniques, due to a certain flexibility of the terms’ use we observed in the literature. In Web of Science, we filtered for the categories of computer science and business
@@
For Search B, illustrated in Fig. 3, we identify 852 candidates for analysis after an initial search and deduplication, as described above. Afterwards, similarly to Search A, we conduct the quality appraisal; after title and abstract screening, 120 articles remain, and 14 after full-test screening. Through a “forward-backward-search”, we identify two additional relevant studies, and two via expert knowledge. Thus, Search B has produced 18 papers we consider relevant for our analysis.
@@
As part of papers added via expert knowledge, we filter out studies in Search A that clearly apply compliance checking techniques instead of conformance checking techniques and add them to Search B as additional expert knowledge; similarly, studies that clearly apply conformance checking techniques instead of compliance checking are screened out and added to Search A as additional expert knowledge. This was done for five studies in total (one study from Search A to B, and four from B to A). For clarity, Figs. 2 and 3 show these studies as part of the “expert knowledge” step, and they have been already included in the preceding paragraphs. After having identified 41/18 (Search A/Search B) relevant studies, we continue with data extraction and synthesis of the extracted information for each of the two searches.
@@ Results
For the action phase, we identify whether expert knowledge is necessary for applying compliance monitoring techniques, whether imperative, declarative, or hybrid techniques [^5] This we determine primarily based on the type of prescriptive model – i.e., declarative, imperative, or hybrid – since the exact technique for regulatory compliance monitoring was not made transparent in all studies we investigated. Nonetheless, the way in which compliance constraints are formalized appears to be a good indicator of the type of technique applied to them. are used (see [carmona2018conformance, de2015alignment]), which process perspectives are assessed (control flow, temporal, data, or resource, see [russellWorkflowPatternsDefinitive2015]), and whether the compliance monitoring techniques are applied manually, in a tool-supported manner, automatically, or in real-time (i.e., online). In contrast to [lyComplianceMonitoringBusiness2015], we also explicitly include the control flow process perspective. However, we do so on a higher level of abstraction than the control flow requirements formulated by [rinderle-maPredictiveComplianceMonitoring2023], which differentiate control flow constraints between existence, absence, and ordering constraints. Subsuming these under the control flow perspective, as in [dunzerConformanceCheckingStateoftheart2019], has been done to describe the application of regulatory compliance monitoring techniques more broadly.
@@ 4 Results
After inductive coding and synthesizing the 59 (41 of Search A, 18 of Search B) relevant studies, we analyse the resulting data, in order to characterize their compliance monitoring mechanisms and their usages. This is done along the three phases of conformance checking projects [carmona2018conformance], being preparation (Sec. 4.1), action (Sec. 4.2), and reflection (Sec, LABEL:sec:results-reflection). We also describe commonalities and differences between the uses found in the two searches. Generally, we can observe, as shown in Fig. 5, that we find far fewer studies that position themselves as compliance checking than those framed as conformance checking studies. We also note that while conformance checking studies were published quite early, they appear to peak slightly later than compliance checking studies, which might hint at a shift in research interest or a change in the positioning of the studies; similarly, compliance checking studies seem to tail off after having reached a peak in 2017.
2009200920102010201120112012201220132013201420142015201520162016201720172018201820192019202020202021202120222022202320232024202402244662202255033331111661166331144331111220112222033111102211110YearStudiesConformanceCompliance
@@ 4.1 Preparation
The set of compliance checking studies, on the other hand, shows a broader distribution of domains (including one study in a healthcare setting): From healthcare, human resources, rent, public services to trade, supply chains and customer service, seven different domains are covered. This seems to suggest a broad applicability of compliance checking in many areas, although we observe a slight focus on healthcare-related studies (five times), trade and supply chain (four times), as well as finance and public services (three times, each).
@@
For compliance checking applications, the situation is slightly: more than half (i.e., ten of 18) of the studies aim to improve compliance checking techniques and their application, one sixth (i.e, three of 18 studies) aim to improve the business process itself, and little over one third (i.e., seven of 18 studies) aims to demonstrate the application of compliance checking. Notably, all three studies that explicitly set out to improve process compliance are situated in the area of trade and supply chains.
@@
Third, we consider how prescriptive models are created in the relevant studies. In all conformance checking and compliance checking studies, we see that no contribution out of the 59 identified studies automatically creates them based on the relevant regulations. In fact, all studies rely on manual input for creation of the required prescriptive model. Notably, the exact operationalization that led to the model is largely not described in detail, and instead, only descriptions or visualizations of them are provided. For example, in [accorsiExploitationProcessMining2012], a BPMN model is created manually with and validated by process experts (in this case, bank managers), and further requirements are provided informally by bank managers and auditors, and manually translated into LTL formulae by the authors.
@@
In contrast to this, only two of the 18 compliance checking studies explicitly mention a validation of the prescriptive model that has been used. [^7] As above, we only noted explicit evaluations, since the remaining seven studies did not mention whether an evaluation took place or not. This may be linked to the prevalent technique improvement goal we see in compliance checking studies: in order to improve compliance checking techniques, the prescriptive models may not need to be evaluated since an evaluated prescriptive model is less relevant here than for demonstrating techniques or improving actual process compliance.
@@
Real-worldSynthetic1010202030304040353566131366OriginStudiesConformanceCompliance
@@
AvailableNot Available1010202030304040883434441414OriginStudies
@@
In the compliance checking studies, we see real-world process event data being used in 13 of 18 studies, and synthetic data in 6 of 18 studies. However, only in four of the real-world studies, this data is actually available – notably, three of them employ BPIC event logs which have been made available for research purposes beforehand [lyComplianceMonitoringBusiness2015, vanbeestCrossInstanceRegulatoryCompliance2023, gongBpCMonRulebasedMonitoring2017].
@@ 4.2 Action
For compliance checking techniques, we observe almost exclusively declarative applications (16 out of 18), regardless of domain. For example, [bukhshComplianceCheckingShipment2017] use ProM’s LTL checker to assess the regulatory compliance of a company’s trade and customs procedures, for which they manually create formal regulatory constraints from relevant statements of the regulations. The prescriptive model is imperative in nature only in two studies, for example in [weidlichEventBasedMonitoring2011] where it serves to derive event queries for use with CEP. The overall prevalence of declarative applications may be linked to mechanisms used for regulatory compliance monitoring, such as CEP, where instead of alignments with an imperative process model, done so in imperative conformance checking techniques [carmona2018conformance], standing queries are evaluated against observed process event data [weidlichEventBasedMonitoring2011] to determine compliance or non-compliance.

### v5

@@ 3. RQ3:
This paper is based on a previous study, where we reviewed conformance checking uses for regulatory compliance monitoring purposes [klessascheckReviewingConformance2024]. Extending our focus, we now additionally investigate usages of further techniques that exist in the research space of regulatory compliance and BPM in our review, and take further aspects such as the origin and availability of process execution data used in the investigated studies into account. We also describe our research method in more detail and, in light of our broadened scope, extend the results and discussion provided herein.
@@ 2.2 Related Work
Building on this framework, Rinderle-Ma et al. [rinderle-maPredictiveComplianceMonitoring2023] classify predictive compliance monitoring approaches along compliance monitoring requirements and capabilities. For doing so, they extend the requirements formulated by Ly et al. [lyComplianceMonitoringBusiness2015] with e.g. additional user requirements or data requirements specific to predictive process compliance monitoring. Similar to Ly et al. [lyComplianceMonitoringBusiness2015], they assess in how far existing literature covers the functionality requirements they elicited, and derive research directions from this assessment. Requirements towards compliance monitoring approaches are elicited from real-world case studies and literature in both studies; however, concrete applications of these approaches are not characterized and investigated for their particular properties. While the requirements are useful for describing the capabilities of compliance monitoring approaches, they are insufficient to characterize how these approaches are applied in a way that goes beyond their ability to meet e.g. modelling, execution, and user requirements. Instead, we believe more emphasis should be placed on how prescriptive models come to be, to what end uses of approaches are reported, and in what ways the results are represented (see also e.g. Rehse et al. [rehse2022process]), in particular for regulatory compliance monitoring.
@@ 3 Research Method
In Search B, we explicitly search for studies that position themselves as applying techniques besides conformance checking (meaning, other compliance checking techniques) for regulatory compliance monitoring. In line with Search A, we include terms commonly associated with regulations. The complete search term ((“compliance checking” OR “compliance monitoring”) AND (“law” OR “guideline” OR “manual” OR “reference” OR “quality” OR “compliance” OR “regulation”)) is applied to the title, abstract, and keyword search, where applicable. [^1] Note that for the search we use compliance monitoring and compliance checking interchangeably for any study that mentions applications of compliance monitoring or checking techniques, due to a certain flexibility of the terms’ use we observed in the literature. In Web of Science, we filtered for the categories of computer science and business.
@@
For Search B, illustrated in Fig. 3, we identify 851 candidates for analysis after an initial search and deduplication, as described above. Afterwards, similarly to Search A, we conduct the quality appraisal; after title and abstract screening, 119 articles remain, and 13 after full-test screening. Through a “forward-backward-search”, we identify two additional relevant studies, and two via expert knowledge. Thus, Search B has produced 17 papers we consider relevant for our analysis.
@@
As part of papers added via expert knowledge, we filter out studies in Search A that clearly apply compliance checking techniques instead of conformance checking techniques and add them to Search B as additional expert knowledge; similarly, studies that clearly apply conformance checking techniques instead of compliance checking are screened out and added to Search A as additional expert knowledge. This was done for five studies in total (one study from Search A to B, and four from B to A). For clarity, Figs. 2 and 3 show these studies as part of the “expert knowledge” step, and they have been already included in the preceding paragraphs. After having identified 41/17 (Search A/Search B) relevant studies, we continue with data extraction and synthesis of the extracted information for each of the two searches.
@@ Results
For the action phase, we identify whether expert knowledge is necessary for applying compliance monitoring techniques, whether imperative, declarative, or hybrid techniques [^5] This we determine primarily based on the type of prescriptive model — i.e., declarative, imperative, or hybrid — since the exact technique for regulatory compliance monitoring was not made transparent in all studies we investigated. Nonetheless, the way in which compliance constraints are formalized appears to be a good indicator of the type of technique applied to them. are used (see [carmona2018conformance, de2015alignment]), which process perspectives are assessed (control flow, temporal, data, or resource, see [russellWorkflowPatternsDefinitive2015]), and whether the compliance monitoring techniques are applied manually, in a tool-supported manner, automatically, or in real-time (i.e., online). In contrast to [lyComplianceMonitoringBusiness2015], we also explicitly include the control flow process perspective. However, we do so on a higher level of abstraction than the control flow requirements formulated by [rinderle-maPredictiveComplianceMonitoring2023], which differentiate control flow constraints between existence, absence, and ordering constraints. Subsuming these under the control flow perspective, as in [dunzerConformanceCheckingStateoftheart2019], has been done to describe the application of regulatory compliance monitoring techniques more broadly.
@@ 4 Results
After inductive coding and synthesizing the 58 (41 of Search A, 17 of Search B) relevant studies, we analyse the resulting data, in order to characterize their compliance monitoring mechanisms and their usages. This is done along the three phases of conformance checking projects [carmona2018conformance], being preparation (Sec. 4.1), action (Sec. 4.2), and reflection (Sec, LABEL:sec:results-reflection). We also describe commonalities and differences between the uses found in the two searches. Generally, we can observe, as shown in Fig. 5, that we find far fewer studies that position themselves as compliance checking than those framed as conformance checking studies. We also note that while conformance checking studies were published quite early, they appear to peak slightly later than compliance checking studies, which might hint at a shift in research interest or a change in the positioning of the studies; similarly, compliance checking studies seem to tail off after having reached a peak in 2017.
2009200920102010201120112012201220132013201420142015201520162016201720172018201820192019202020202021202120222022202320232024202402244662202255033331111661166331144331111220112222033111101111110YearStudiesConformanceCompliance
@@ 4.1 Preparation
The set of compliance checking studies, on the other hand, shows a broader distribution of domains (including one study in a healthcare setting): From healthcare, human resources, rent, public services to trade, supply chains and customer service, seven different domains are covered. This seems to suggest a broad applicability of compliance checking in many areas, although we observe a slight focus on healthcare-related studies (five times), trade and supply chain (four times), as well as finance and public services (three times, two times).
@@
For compliance checking applications, the situation is slightly: more than half (i.e., ten of 17) of the studies aim to improve compliance checking techniques and their application, only two of 17 studies aim to improve the business process itself, and little over two fifths (i.e., seven of 17 studies) aim to demonstrate the application of compliance checking. Notably, all three studies that explicitly set out to improve process compliance are situated in the area of trade and supply chains.
@@
Third, we consider how prescriptive models are created in the relevant studies. In all conformance checking and compliance checking studies, we see that no contribution out of the 58 identified studies automatically creates them based on the relevant regulations. In fact, all studies rely on manual input for creation of the required prescriptive model. Notably, the exact operationalization that led to the model is largely not described in detail, and instead, only descriptions or visualizations of them are provided. For example, in [accorsiExploitationProcessMining2012], a BPMN model is created manually with and validated by process experts (in this case, bank managers), and further requirements are provided informally by bank managers and auditors, and manually translated into LTL formulae by the authors.
@@
In contrast to this, only two of the 17 compliance checking studies explicitly mention a validation of the prescriptive model that has been used. [^7] As above, we only noted explicit evaluations, since the remaining seven studies did not mention whether an evaluation took place or not. This may be linked to the prevalent technique improvement goal we see in compliance checking studies: in order to improve compliance checking techniques, the prescriptive models may not need to be evaluated since an evaluated prescriptive model is less relevant here than for demonstrating techniques or improving actual process compliance.
@@
Real-worldSynthetic1010202030304040353566121266OriginStudiesConformanceCompliance
@@
AvailableNot Available1010202030304040883434441313OriginStudies
@@
In the compliance checking studies, we see real-world process event data being used in twelve of 17 studies, and synthetic data in six of 17 studies. However, only in four of the real-world studies, this data is actually available – notably, three of them employ BPIC event logs which have been made available for research purposes beforehand [lyComplianceMonitoringBusiness2015, vanbeestCrossInstanceRegulatoryCompliance2023, gongBpCMonRulebasedMonitoring2017].
@@ 4.2 Action
For compliance checking techniques, we observe almost exclusively declarative applications (15 out of 17), regardless of domain. For example, [bukhshComplianceCheckingShipment2017] use ProM’s LTL checker to assess the regulatory compliance of a company’s trade and customs procedures, for which they manually create formal regulatory constraints from relevant statements of the regulations. The prescriptive model is imperative in nature only in two studies, for example in [weidlichEventBasedMonitoring2011] where it serves to derive event queries for use with CEP. The overall prevalence of declarative applications may be linked to mechanisms used for regulatory compliance monitoring, such as CEP, where instead of alignments with an imperative process model, done so in imperative conformance checking techniques [carmona2018conformance], standing queries are evaluated against observed process event data [weidlichEventBasedMonitoring2011] to determine compliance or non-compliance.